# Cosmological Consequences of Varying Couplings in Gravity Action

### v1

@@ Abstract
We develop a Lagrangian formulation for gravity with matter where the gravitational couplings are universally treated as being field-dependent. The solutions for FLRW geometries and the associated time evolution of the Newton and cosmological couplings are found. The distance-redshift relations are shown to prefer a slowly growing Newton’s coupling alongwith a negative equation of state ($w\geq-\frac{1}{3}$) for the matter fluid at the present epoch of accelerated expansion, while ruling out Dirac’s large number hypothesis. We obtain an improved bound on $\dot{G}(t)$ as: $1.67\times 10^{-11}yr^{-1}<\frac{\dot{G}}{G}<3.34\times 10^{-11}yr^{-1}$ in the context of supernova cosmology, as well as a new constraint on $\dot{\Lambda}(t)$ as: $-0.67\times 10^{-11}yr^{-1}<\frac{\dot{\Lambda}}{\Lambda}<-1.34\times 10^{-11}yr^{-1}$. Based on this formulation, we also present a dynamical solution to the ‘cosmic coincidence’ problem.
@@ I Introduction
The question as to whether the couplings in the gravity action are constants or are varying in time is one of fundamental importance. Early attempts exploring the idea of a time-dependent Newton’s coupling could be traced back to Dirac’s large number hypothesis [1, 2], where the precise form of the variation was invoked towards explaining the origin of large dimensionless numbers ($\sim 10^{39}$) in nature. Later on, Brans and Dicke had proposed a gravitational theory with varying $G$ from a Lagrangian perspective [3] with an aim to connect gravity to Mach’s principle. The possibility of varying couplings have since been considered from various perspectives [4, 5, 6, 7, 8, *smolin1]. Phenomenological models of running couplings have also been richly investigated [10, 11, 12, 13].
However, the problem of formulating a universal Lagrangian framework where all possible gravitational couplings could vary (in time) requires a fresh perspective. Here, we set up a Lagrangian formulation for gravity where the possible time-evolution of both Newton’s and cosmological coupling ($G$ and $\Lambda$, respectively) could be described through a single scalar field without introducing any modification (or, any potential) in the matter sector. Based on the resulting dynamics, we analyze the possible cosmological consequences in detail. The resulting theory is not the same as the one where the couplings are treated as constants from the outset, namely, before varying the action. The cosmological implications found here goes beyond the scope of standard FLRW cosmology and the $\Lambda CDM$ model, laying down rich future prospects.
@@ III FLRW cosmology: Field equations
The set of three equations above specifies the complete system (for a given matter field), which is underdetermined. Here we would solve the system by assuming a power law and an equation of state $w\equiv\frac{P}{\rho}=const.$
@@ IV.1 Generic features
a. For $w<\frac{1}{3}$, positivity of density requires: $\frac{(\alpha+2)\lambda}{\alpha}<0$. Hence, for a positive $\lambda$ the Newton’s coupling $\xi^{-1}\equiv G$ must grow in time: $-2<\alpha<0$. For $\lambda<0$, it could be either decaying ($20<\alpha<$ 2) or growing ($\alpha<-2$), corresponding to a slower and faster variation, respectively. The ultra-relativistic equations of state $w>\frac{1}{3}$, on the other hand, exhibit exactly the opposite behaviour.
@@
d. An accelerated expansion of the universe requires $\frac{2-\alpha}{3(1+w)}>1$, which implies that the cosmological coupling $\lambda$ must be positive. Thus, for both $w>\frac{1}{3}$ and $w<\frac{1}{3}$, acceleration is possible. These two limits are relevant for the early and late universes, respectively.
@@ A. Stiff fluid:
For $P=\rho$, $a\sim t^{\frac{2-\alpha}{6}},~{}\rho\sim a^{-6}$. The two classes of solutions are defined by $\lambda>0$, associated with either decaying $G(t)$ ($02>\alpha>$ 0) or growing $G(t)$ ($\alpha<-2$) and $\lambda<0$. Thus, for a relatively slow variation of the coupling, the stiff phase which could dominate only at the early universe (expanding) is expected to exhibit a decaying $G(t)$.
@@ B. Dust:
For $P=0$, $a\sim t^{\frac{2-\alpha}{3}},~{}\rho\sim a^{-3}$. Note that for positive $\lambda$, solutions now correspond to only growing $G$ ($-2<\alpha<0$), whereas negative $\lambda$ is associated with either $02>\alpha>$ 0 (decaying) or $\alpha<-2$ (growing). Remarkably, the universe could accelerate even in the matter dominated era for positive $\lambda$ provided $\alpha<-1$. Assuming a positive $\lambda$, the age of a matter-dominated universe is given by: $\frac{2}{3H_{0}}\leq t_{0}<\frac{4}{3H_{0}}$.
@@ V Solutions for radiation matter
$\lambda(\phi)~{}=~{}C\phi^{2},~{}C~{}\equiv~{}\mathrm{const.}$ (12)
@@
[…]
@@
a. As in the previous case, these solutions also fall into two classes: decaying $G(t)$ for $20<\alpha<$ 2 and growing $G(t)$ for $\alpha<0$.
@@
c. For the age of a radiation dominated universe, we have $t_{0}\leq\frac{1}{2H_{0}}$ for a decaying $G(t)$ and $t_{0}<\frac{1}{H_{0}}$ for a growing $G(t)$.
@@ VI Distance-Redshift relation
[…]
@@
Figure 1: Comoving distance for dust domination (the null curve corresponds to $\alpha=0$, encoding pure matter-domination in a $\Lambda CDM$ Universe)
Figure 2: Comoving distance for a negative EOS
@@ VII Constraints on the variation of $G$ and $\Lambda$
The rate of variation of the Newton’s coupling at present reads: $\frac{\dot{G}}{G}=-\frac{\dot{\xi}}{\xi}=-\frac{\alpha}{t_{0}}$. Using the upper and lower bounds on the age, we find a dimensionless constraint as: $-\frac{\alpha H_{0}}{2}<\frac{\dot{G}}{G}<-\alpha H_{0}$. Assuming $w=-\frac{1}{3}$ at the present epoch and $\alpha\approx-\frac{1}{2}$ using the distance-redshift diagram, we find:
@@
where $H_{0}=67$ $kms^{-1}Mpc^{-1}$. Note that this bound is stronger than the existing ones derived within supernova cosmology assuming a $\Lambda CDM$ model and extending upto about $z\sim 1$ [18, 19].
@@
Unlike $G$, this constraint is independent of $\alpha$. This numerically translates to: $-6.7\times 10^{-11}yr^{-1}>\frac{\dot{\Lambda}}{\Lambda}>-13.4\times 10^{-11}yr^{-1}$.
@@ VIII A solution to the Cosmic Coincidence Problem
Given the broad range of the current limits on $\frac{\dot{G}}{G}$ presented within various contexts [11, *uzan, 20], it might be premature to decide whether the exponent of variation $|\alpha|$ lies closer to 00 or is of order unity. Hence, let us consider both possibilities as below.
As confirmed by observations, let us assume that the universe has entered a phase of accelerated expansion after the epoch of matter domination ($\lambda>0,~{}w\lesssim 0$). Given the inequality $\Omega_{m}>\Omega_{\lambda}$, we have: $\alpha>-\frac{4}{3}$ for a lower bound of the exponent during matter domination. Hence, we may assume that $\alpha$ is close to $-1$ ($>-1$ for decelerating universe) during this epoch, for which the ratio becomes: $\frac{\Omega_{m}}{\Omega_{\lambda}}|_{w=0}\sim 2$.
Consistency of the $a_{0}H_{0}\chi(z)$ vs $z$ diagram (figure $2$) with the supernovae data, however, indicates that for a slower variation of $\alpha$ than above ($-1<\alpha<0$), an accelerated universe now is consistent with a negative EOS $0>w\geq-\frac{1}{3}$. The corresponding ratio then becomes: $\frac{\Omega_{m}}{\Omega_{\lambda}}|_{w=-\frac{1}{3}}=|-\frac{\alpha+2}{\alpha}|=\frac{1}{2}\frac{\Omega_{m}}{\Omega_{\lambda}}|_{w=0}$. Thus, for a given $\alpha$, the density ratio remains of a similar order of magnitude between the current and matter dominated epoch, which define most of the lifetime of the universe.
Finally, let us also consider the question if the ratio could be widely different in order of magnitude during the early stages of evolution. Since $w>\frac{1}{3}$ is expected dominate the density at early times in an expanding universe, we may take the stiff EOS as an extreme representative of the early phase. A positive $\lambda$ and a relatively slowly varying $G(t)$ implies $20<\alpha<$ 2. In this epoch we obtain: ${\frac{\Omega_{m}}{\Omega_{\lambda}}}=\frac{\alpha+2}{\alpha}$, which, again, is of a similar order of magnitude as in the current and matter dominated epochs irrespective of whether $|\alpha|$ is closer to 0 or 1.
@@ IX Concluding remarks
There are a few significant imports of our results here. The effective $\Lambda$, superceding the ‘dark energy’ component, becomes directly related to the energy momentum tensor of the matter fluid through the field equations. The radiation dominated epoch emerges as a critical phase where $\Lambda$ vanishes. During any other epoch ($w\neq\frac{1}{3}$), the matter and $\Lambda$ densities exhibit a constant ratio, in contrast with the standard $\Lambda$CDM model based on constant couplings. This reflects a new method to solve the cosmic coincidence problem involving these two densities dynamically, and could supercede the quintessance models or their variants [4, 14]. In principle, our formulation admits the possibility of an accelerated expansion both in the early and late universe. These features show that the scope of the analysis here goes beyond standard FLRW or the scalar-field based cosmology, and could be used to address some of the outstanding problems in cosmology.
It is intriguing to observe that $\xi\sim t$ emerges as a particular solution of the general class, resembling the time-evolution of $G$ as conjectured in Dirac’s large number hypothesis (LNH). However, as a theoretical framework, the critical differences between our formulation and the latter are obvious. To emphasize, the action principle here describes the time evolution of both $G$ and $\Lambda$ while using a single metric (the generalization to include any other gravitational coupling is straightforward), and leads to no exotic feature such as a spontaneous creation of matter in empty space or otherwise, in contrast with the LNH.
@@ Acknowledgements.
This work is supported by the MATRICS grant MTR/2021/000008, SERB, Govt. of India.

### v2

@@ Abstract
We develop a Lagrangian formulation for gravity with matter where the gravitational couplings are universally treated as being field-dependent. The solutions for FLRW geometries and the associated time evolution of the Newton and cosmological couplings are found. The distance-redshift relations are shown to prefer a slowly growing Newton’s coupling alongwith a negative equation of state ($w\geq-\frac{1}{3}$) for the matter fluid at the present epoch of accelerated expansion, while ruling out the evolution of $G$ predicted by Dirac’s large number hypothesis. We obtain an improved bound on $\dot{G}(t)$ as: $1.67\times 10^{-11}yr^{-1}<\frac{\dot{G}}{G}<3.34\times 10^{-11}yr^{-1}$ in the context of supernova cosmology, as well as a new constraint on $\dot{\Lambda}(t)$ as: $0.67\times 10^{-11}yr^{-1}<\frac{\dot{\Lambda}}{\Lambda}<-1.34\times 10^{-11}yr^{-1}$. Based on this formulation, we also present a dynamical solution to the ‘cosmic coincidence’ problem.
@@ I Introduction
The question as to whether the couplings in the gravity action are constants or are varying in time is one of fundamental importance. Early attempts exploring the idea of a time-dependent Newton’s coupling could be traced back to Dirac’s large number hypothesis [1, 2], where the precise form of the variation was invoked towards explaining the origin of large dimensionless numbers ($\sim 10^{39}$) in nature. Later on, Brans and Dicke had proposed a gravitational action underlying a variable $G$ [3] with an aim to connect gravity to Mach’s principle. The possibility of varying couplings have since been considered from various perspectives [4, 5, 6, 7, 8, *smolin1]. Phenomenological models of running couplings have also been richly investigated [10, 11, *uzan, 13].
However, the problem of formulating a universal Lagrangian framework where all possible gravitational couplings could vary (in time) requires a fresh outlook. Here, we set up a Lagrangian formulation for gravity where the possible time-evolution of both Newton’s and cosmological coupling ($G$ and $\Lambda$, respectively) could be described through a single scalar field without introducing any modification (or, any potential) in the matter sector. Based on the resulting dynamics, we analyze the possible cosmological consequences in detail. The resulting theory is not the same as the one where the couplings are treated as constants from the outset, namely, before varying the action. The cosmological implications found here goes beyond the scope of standard FLRW cosmology and the $\Lambda CDM$ model, laying down rich future prospects.
@@ III FLRW cosmology: Field equations
The set of three equations above specifies the complete system (for a given matter field), which is underdetermined. In order to solve for $a(t),\xi(t),\lambda(t),\rho(t),P(t)$, we need two additional conditions. Here we would solve the system by assuming a power law ansatz and a constant equation of state $w\equiv\frac{P}{\rho}=const.$ Since the standard FLRW solutions in Einstein theory with a constant $w$ emerge as power law, this makes a direct comparison of the resulting dynamics in the two cases possible.
@@ IV.1 Generic features
a. For $w<\frac{1}{3}$, positivity of density requires: $\frac{(\alpha+2)\lambda}{\alpha}<0$. Hence, for a positive $\lambda$ the Newton’s coupling $\xi^{-1}\equiv G$ must grow in time: $2<\alpha<0$. For $\lambda<0$, it could be either decaying ($0<\alpha<2$) or growing ($\alpha<-2$), corresponding to a slower and faster variation, respectively. The ultra-relativistic equations of state $w>\frac{1}{3}$, on the other hand, exhibit exactly the opposite behaviour.
@@
d. An accelerated expansion of the universe requires $\frac{2-\alpha}{3(1+w)}>1$. Thus, for both $w>\frac{1}{3}$ and $w<\frac{1}{3}$, acceleration is possible. These two limits are relevant for the early and late universes, respectively.
@@ A. Stiff fluid:
For $P=\rho$, $a\sim t^{\frac{2-\alpha}{6}},~\rho\sim a^{-6}$. The two classes of solutions are defined by $\lambda>0$, associated with a slowly decaying ($2>\alpha>0$) or a rapidly growing ($\alpha<-2$) $G(t)$, and $\lambda<0$ with a slowly growing ($2<\alpha<0$) $G(t)$. This stiff phase could dominate only in the early universe.
@@ B. Dust:
For $P=0$, $a\sim t^{\frac{2-\alpha}{3}},~\rho\sim a^{-3}$. Note that for positive $\lambda$, solutions now correspond to only growing $G$ ($2<\alpha<0$), whereas negative $\lambda$ is associated with either $2>\alpha>0$ (decaying) or $\alpha<-2$ (growing). Remarkably, the universe could accelerate even in the matter dominated era for positive $\lambda$ provided $\alpha<-1$. Assuming a positive $\lambda$, the age of a matter-dominated universe is given by: $\frac{2}{3H_{0}}\leq t_{0}<\frac{4}{3H_{0}}$.
@@ V Solutions for radiation matter
$\lambda(\xi)~=~C\xi^{2},~C~\equiv~\mathrm{const.}$ (12)
@@
$\frac{\lambda}{\xi}$ $~=$~ $\frac{\lambda_{0}}{\xi_{0}}\Big{(}\frac{t}{t_{0}}\Big{)}^{\alpha}~=~\frac{3}{2}\Big{[}-\partial_{t}\big{(}\frac{\dot{a}}{a}\big{)}+\frac{3\alpha\dot{a}}{2ta}+\frac{\alpha}{4t^{2}}(\alpha-2)\Big{]}$
$\frac{\rho}{\xi}$ $~=$~ $\frac{3}{2}\Big{[}-\partial_{t}\big{(}\frac{\dot{a}}{a}\big{)}+\frac{\alpha\dot{a}}{2ta}+\frac{\alpha}{4t^{2}}(\alpha+2)\Big{]}$
@@
a. As in the previous case, these solutions also fall into two classes: decaying $G(t)$ for $0<\alpha<2$ and growing $G(t)$ for $\alpha<0$.
@@
c. For the age of a radiation dominated universe, we have $t_{0}\leq\frac{1}{2H_{0}}$ for a decaying $G(t)$ and $t_{0}>\frac{1}{2H_{0}}$ for a growing $G(t)$.
@@ VI Distance-Redshift relation
$\chi(z)$ $~=$~ $\frac{1}{a_{0}}\int_{0}^{z}\frac{dz}{H(z)}~=~\frac{1}{H_{0}}\Bigg{[}\frac{(1+z)^{1-\frac{3(1+w)}{2-\alpha}}-1}{1-\frac{3(1+w)}{2-\alpha}}\Bigg{]}$ (14)
$~=$~ $\frac{1}{a_{0}H_{0}}\Big{[}z~-~\frac{3(1+w)}{2(2-\alpha)}z^{2}~+~\Bigg{(}\frac{1+w}{2(2-\alpha)}\Bigg{)}\Bigg{(}\frac{3(1+w)}{(2-\alpha)}+1\Bigg{)}z^{3}~+~o(z^{4})\Bigg{]}$
@@
Figure 1: Comoving distance for dust matter domination ($w=0$) for different positive and negative exponents $\alpha$ ($1,\frac{1}{2},0,-\frac{1}{2},-\frac{3}{2}$). The null curve ($\alpha=0$) corresponds to a pure matter-domination ($\Omega_{\Lambda}=0$) in a $\Lambda CDM$ Universe. The thick black line refers to the $\Lambda CDM$ curve assuming $\Omega_{m}=0.32,\Omega_{\Lambda}=0.68$.
Figure 2: Comoving distance for the EOS $w=-\frac{1}{3}$, plotted for different positive and negative exponents $\alpha$ ($1,\frac{1}{2},-\frac{1}{2},-1$). $\alpha=1$ corresponds to Dirac’s conjecture (LNH) $G\sim t^{-1}$, which is clearly disfavoured.
@@ VII Constraints on the variation of $G$ and $\Lambda$
The rate of variation of the Newton’s coupling at present reads: $\frac{\dot{G}}{G}=-\frac{\dot{\xi}}{\xi}=-\frac{\alpha}{t_{0}}$. Using the upper and lower bounds on the age, we find a dimensionless constraint as: $\frac{\alpha H_{0}}{2}<\frac{\dot{G}}{G}<-\alpha H_{0}$. Assuming $w=-\frac{1}{3}$ at the present epoch and $\alpha\approx-\frac{1}{2}$ using the distance-redshift diagram, we find:
@@
where $H_{0}=67$ $kms^{-1}Mpc^{-1}$. Note that this bound is stronger than the existing ones obtained within supernova cosmology assuming a $\Lambda CDM$ model and extending upto about $z\sim 1$ [18, 19].
@@
Unlike $G$, this constraint is independent of $\alpha$. This numerically translates to: $6.7\times 10^{-11}yr^{-1}>\frac{\dot{\Lambda}}{\Lambda}>-13.4\times 10^{-11}yr^{-1}$.
@@ VIII A solution to the Cosmic Coincidence Problem
Given the broad range of the current limits on $\frac{\dot{G}}{G}$ presented within various contexts [11, *uzan, 20], it might be premature to decide whether the exponent of variation $|\alpha|$ lies closer to 0 or is of order unity. Hence, let us consider both possibilities as below.
As confirmed by observations, let us assume that the universe has entered a phase of accelerated expansion after the epoch of matter domination ($\lambda>0,~w\lesssim 0$). Given the inequality $\Omega_{m}>\Omega_{\lambda}$, we have: $\alpha>-\frac{4}{3}$ for a lower bound of the exponent during matter domination. Hence, we may assume that $\alpha$ is close to 1-1 (>1>-1 for decelerating universe) during this epoch, for which the ratio becomes: $\frac{\Omega_{m}}{\Omega_{\lambda}}|_{w=0}\sim 2$.
Consistency of the $a_{0}H_{0}\chi(z)$ vs $z$ diagram (figure $2$) with the supernovae data, however, indicates that for a slower variation of $\alpha$ than above ($1<\alpha<0$), an accelerated universe now is consistent with a negative EOS $0>w\geq-\frac{1}{3}$. The corresponding ratio then becomes: $\frac{\Omega_{m}}{\Omega_{\lambda}}|_{w=-\frac{1}{3}}=|-\frac{\alpha+2}{\alpha}|=\frac{1}{2}\frac{\Omega_{m}}{\Omega_{\lambda}}|_{w=0}$. Thus, for a given $\alpha$, the density ratio remains of a similar order of magnitude between the current and matter dominated epoch, which define most of the lifetime of the universe.
Finally, let us also consider the question if the ratio could be widely different in order of magnitude during the early stages of evolution. Since $w>\frac{1}{3}$ is expected to dominate the density at early times in an expanding universe, we may take the stiff EOS as an extreme representative of the early phase. A positive $\lambda$ and a relatively slowly varying $G(t)$ implies $0<\alpha<2$. In this epoch we obtain: ${\frac{\Omega_{m}}{\Omega_{\lambda}}}=\frac{\alpha+2}{\alpha}$, which, again, is of a similar order of magnitude as in the current and matter dominated epochs irrespective of whether $|\alpha|$ is closer to 0 or 1.
@@ IX Concluding remarks
[…]
It is intriguing to observe that $\xi\sim t$ emerges as a particular solution of the general class, resembling the time-evolution of $G$ as conjectured in Dirac’s large number hypothesis (LNH). However, the critical differences between our formulation and the latter are obvious. To emphasize, the action principle here describes the time evolution of both $G$ and $\Lambda$ while being based on a single metric (the generalization to include any other gravitational coupling is straightforward), and leads to no exotic feature such as a spontaneous creation of matter in empty space, in contrast with the LNH.
@@ Acknowledgements.
The support of the MATRICS grant MTR/2021/000008, SERB (ANRF), Govt. of India is gratefully acknowledged.